OpenStudy (anonymous):

Write an equation of a vertical line passing through the point (6, 3).

OpenStudy (anonymous):

what doesn't change for a vertical line?

OpenStudy (anonymous):

or, what is another point on the line?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

i asked a couple of questions (see previous responses)... do you have any answers or comments?

OpenStudy (anonymous):

sorry im new to this site bro

OpenStudy (anonymous):

and stupid at math lol

OpenStudy (anonymous):

Slope-intercept form: y = mx + b Point-slope form: y-y1= m(x-x1) Standard form: Ax + By = C; where A is a positive value

OpenStudy (anonymous):

m=change in y=y2-y1 change in x=x2=x1

OpenStudy (anonymous):

looks like you've told yourself that enough times that you're actually starting to believe it. what a shame. draw a vertical line going through the given point and find another point on that line

OpenStudy (anonymous):